# Hypergeometric Distributions and joint families of elliptic curves Brian Grove and Hasan Saad

### v1

@@ Abstract.
Gaussian hypergeometric functions, Sato–Tate type distributions, Elliptic curves
2020 Mathematics Subject Classification:
11G20, 11T24, 33E50
1. Introduction and Statement of Results
@@
The limiting distributions of these finite field hypergeometric functions have been studied in multiple works [OSS23, Saa23, OPSS24, Gro24]. To make this precise, we first recall the definition of these finite field hypergeometric functions. Let be a prime, a fixed primitive th root of unity, and let be a generator of the character group of Furthermore, let and be two collections of rational numbers and let be the greatest common denominator of 111We say that the hypergeometric datum is of length In this notation, if and then we have [McC12b] the finite field hypergeometric function222The function is a variant of the hypergeometric functions originally defined by Greene.
@@
Here we determine the limiting behavior of for certain hypergeometric data of length four for which the condition can be removed (see Theorem 1.3 of [BCM15]). The following theorem gives the limiting moments.
@@
The Catalan numbers are ubiquitous in mathematics ([Sta15]). In relation to our results, we have that these Catalan numbers
arise (see (5.3) of [FKRS12]) as the even moments of the traces of the Lie group More precisely, if then we have
@@
We give the following remarks regarding Theorems 1.4, 1.7, 1.8, and Corollary 1.5.
@@
We remark that Theorem 1.7 follows immediately from the work of Michel [Mic95]. We add it here for completeness and to illustrate the connection between the geometric methods used here and the modular methods used in [OSS23, OPSS24, Gro24].
@@
Geometrically, the arithmetic of the functions in Theorem 1.8 is related to twisted symmetric squares of the Clausen family of elliptic curves (see Lemma 2.2).
The moments in Theorem 1.8 arise (see [PV04]) as the even moments of the traces of the Lie group More precisely, if is even, then we have
@@
This paper is organized as follows. In Section 2, we recall important facts on the relations between the functions mentioned above and families of elliptic curves. In Section 3, we recall necessary background from the theory of étale cohomology and important facts on the first cohomology of elliptic curves. In Section 4, we recall facts from representation theory. In Section 5, we apply the results from étale cohomology and representation theory to compute the mixed moments of families of elliptic curves under a certain “independence” condition of their first cohomologies. In Section 6, we describe this “independence” condition in terms of isogenies of families of elliptic curves viewed as elliptic curves over In Section 7, we recall the method of moments and derive the necessary distributions for Corollaries 1.3 and 1.5. In Section 8, we conclude with the proofs of Theorems 1.1 through 1.8 and Corollaries 1.3 and 1.5.
@@ Acknowledgements
The authors would like to thank Ling Long and Fang-Ting Tu for the numerous discussions and advice which highly contributed to the quality of this paper.
The authors would also like to thank Bill Hoffman and Ken Ono for their multiple comments, discussions, and helpful remarks. The authors also thank Wanlin Li who pointed out an error in an earlier draft. The first author was partially supported by a research assistantship from the Louisiana State University (LSU) Department of Mathematics.
@@ 2. The functions and the arithmetic of elliptic curves
Proof.
If is not a square, the fact that follows immediately from Theorem 2.2 and Lemma 2.6 of [MT24].
To prove (2.1) when note that Corollary 2.3 and Lemma 2.6 of [MT24] imply that
In order to prove Theorems 1.1 through 1.8, we describe the traces of an elliptic curve as traces of Frobenius morphisms on an -adic étale sheaf. This perspective allows us to use the resuls and heavy machinery of étale cohomology which we assume the reader is familiar with. In Subsection 3.1, for the reader’s convenience, we recall the necessary background from the theory of étale cohomology. In Subsection 3.2 we recall some background on the sheaves which encode finite-field point counts on families of elliptic curves. Those sheaves “are” essentially the rational -adic Tate modules when the family is viewed as an elliptic curve defined over the function field (see Section 6).
@@
For background on the applications of étale cohomology in analytic number theory, the reader is referred to [FKMS19], [FKM14], [FKM15], [Kat90b]. For a more geometric perspective, the reader is referred to [Mil80].
@@
In order to prove Theorems 1.1 and 1.4, we will describe the moments of the values as sums of the form where is an appropriately chosen étale sheaf. To determine the asymptotics of these moments, we express the sums over as sums of traces of acting on cohomology groups associated to the sheaf
@@
[FKMS19, Section 3.4 and Example 4.2.2]
Let be a prime a character on and Then there exists a rank- -adic sheaf , pure of weight , lisse on and tamely ramified at and such that
@@
Parts (1) and (2) are Lemma 3.1 of [Mic95]. To see part (3), note that the smooth base change theorem (see Chapter VI, Corollary 4.2 of [Mil80]) implies that is isomorphic as a Galois representation to the where is viewed as an elliptic curve over The self-duality then follows by Poincaré duality (see Chapter VI, Corollary 11.2 of [Mil80]) which in this case is nothing but the Weil pairing.
@@
Lemma 3.10(p. 135 of [Mic95], Section 4 of [Del80]).
Let such that the fiber has multiplicative reduction. Then, in some basis, the representation restricted to is given by
where is the canonical surjection from to
@@ 4. Background from Representation Theory
Here we recall facts from representation theory which allow for computation of moments using the representations described in Section 3. Aside from the technical details, this section is essentially a generalization of some of the tools used in [OSS23, OPSS24, Gro24]. More precisely, the results in this section combined with Lemma 3.9 elucidate why the Chebyshev polynomial sums
are precisely the sums with small asymptotic magnitude. Furthermore, this section elucidates why the Catalan numbers are the moments and transforms the induction arguments on Chebyshev polynomials into direct explicit formulas.
@@
In order to compute the moments in Theorems 1.1 and 1.7, we decompose the tensor product representations into a direct sum of symmetric power representations and a trivial representation. The aim of this is to isolate the irreducible nontrivial components, which by Lemma 3.4, Theorem 3.2, and Theorem 3.3 have small contributions to the moments. The following well-known lemma (for example, see 2.2. of [LZ10]) provides the required direct sum decomposition.
@@
If then this is trivial. Now suppose
Using Lemma 4.3, we have that
@@
Putting all this together, we have our claim.
@@
In order to use the geometric monodromy computation of Deligne in Lemma 3.9, we connect the irreducibility of with
Recall that the geometric mondromy group is the Zariski closure of in The following well-known lemma (for example, see Exercise 7.1.6 of [Kow14]) relates the irreducibility of a representation
@@
Here we write the proof for the reader’s convenience.
If is irreducible, it follows immediately that acts irreducibly. Now, suppose the standard representation of is irreducible, and let be a subspace of such that
@@
It is clear that is a Zariski closed algebraic subgroup of which contains This implies that Since acts irreducibly, or proving our claim.
@@
Let be the basis in the statement of the lemma, and let be an element of If the subspace generated by is invariant under the action of a direct computation shows that
@@ 5. Moments of Frobenius Traces
Here we assemble the results from Sections 3 and 4 to compute the moments and mixed-moments that are required for Theorems 1.1 through 1.8.
@@
where the representations are irreducible pairwise nonisomorphic nontrivial representations of and where is the trivial representation. Part (1) of Lemma 3.9 states that the Zariski closure of Therefore, by Lemma 4.7, we have that the are precisely the symmetric power representations and are the multiplicities in Lemma 4.3.
Now, note that by Theorem 3.6, we have that the Zariski closure of is given by Therefore, Lemma 4.8 implies that for we have that
is an irreducible representation of It is also clear that and are irreducible.
Therefore, the multiplicity of the trivial representation in the decomposition of is equal to and then Lemma 4.3 shows our claim.
@@
If is an irreducible representation as in the statement, it is clear that viewed as an -adic sheaf, is lisse on Therefore, Theorem 3.2 gives us that
@@
by identifying the sum with the trace of Frobenius on the cohomology group where
Now, for let and let be the family defined in Section 2. In this case, we have an explicit description of the Frobenius traces on this “mysterious” cohomology group in terms of the traces of Hecke operators on spaces of cusp frorms. For example, if and is even, then (see Section 4.2 of [Gro24]) the trace of the Frobenius map is given by
@@
where is the multiplicity of the trivial representation in Combining this with Lemma 5.2, we obtain our claim.
@@
Now, note that these irreducible representations are symmetric powers of By Lemma 4.6, we have that is irreducible. Moreover, it is clear that this representation is never trivial. Therefore, the same argument as in the proof of Lemma 5.3 shows our claim.
@@
Just as in Proposition 5.1, the cohomology groups that show up in the proof of Proposition 5.5 have explicit traces using modular forms. For more details, see Proposition 2.2 and the proof of Proposition 6.3 in [Saa23].
@@
which proves our claim.
@@ 6. Tate Modules and Isogenies
First, by Lemma 3.10 and our assumption of good or multiplicative reduction, it is clear that and must have the same points of ramification. By Lemma 4.9, we have that must be lisse on Furthermore, by Lemma 3.9, there exists such that if , then and are tamely ramified at and therefore, is tamely ramified at Therefore, is geometrically constant (see Example 5.2 (f) of Chapter I of [Mil80]).
@@
Now, take to be a nonsquare in and consider By comparing the traces of Frobenius of and the quasi-orthogonality relations of irreducible sheaves (see Theorem 5.2 of [FKMS19]) imply that, up to enlarging we have for all primes
This immediately follows from Lemmas 6.1, 6.4, 6.5, and Theorem 6.2.
@@ 7. Some Distributions
To obtain Corollaries 1.3 and 1.5, we combine Theorems 1.1 and 1.4 with the method of moments. In this section, we compute the required probability density functions associated to a given sequence of moments or mixed moments.
@@
and our claim follows immediately.
@@
has multiplicative reduction at whereas has good reduction at Therefore, Proposition 5.1 applied to combined with Lemma 5.7 gives the claimed moments.
This follows from Lemma 2.2 and Proposition 5.5.

### v2

@@ Abstract.
Gaussian hypergeometric functions, Sato–Tate type distributions, Elliptic curves2020 Mathematics Subject Classification:
11G20, 11T24, 33E501. Introduction and Statement of Results
@@
The limiting distributions of these finite field hypergeometric functions have been studied in multiple works [OSS23, Saa23, OPSS24, Gro24]. To make this precise, we first recall the definition of these finite field hypergeometric functions. Let be a prime, a fixed primitive th root of unity, and let be a generator of the character group of Furthermore, let and be two collections of rational numbers and let be the least common multiple of the denominators of 111We say that the hypergeometric datum is of length In this notation, if and then we have [McC12b] the finite field hypergeometric function222The function is a variant of the hypergeometric functions originally defined by Greene.
@@
Here we determine the limiting behavior of for certain hypergeometric data of length four.
Remark 1.1.
For the hypergeometric data we consider, the condition can be weakened to (see Theorem 1.3 of [BCM15]). Therefore, since we are concerned with the asymptotic behavior as in what follows, we have no restrictions on
The following theorem gives the limiting moments.
@@
The Catalan numbers are ubiquitous in mathematics ([Sta15]). In relation to our results, we have that these Catalan numbers arise (see (5.3) of [FKRS12]) as the even moments of the traces of the Lie group More precisely, if then we have
@@
- (2)
@@
- (4)
- (5)
@@
This paper is organized as follows. In Section 2, we recall important facts on the relations between the functions mentioned above and families of elliptic curves. In Section 3, we recall necessary background from the theory of étale cohomology and important facts on the first cohomology of elliptic curves. In Section 4, we recall facts from representation theory. In Section 5, we apply the results from étale cohomology and representation theory to compute the mixed moments of families of elliptic curves under a certain “independence” condition of their first cohomologies. In Section 6, we describe this “independence” condition in terms of isogenies of families of elliptic curves viewed as elliptic curves over In Section 7, we recall the method of moments and derive the necessary distributions for Corollaries 1.4 and 1.6. In Section 8, we conclude with the proofs of Theorems 1.2 through 1.9 and Corollaries 1.4 and 1.6.
@@ Acknowledgements
The authors would like to thank Ling Long and Fang-Ting Tu for the numerous discussions and advice which highly contributed to the quality of this paper. The authors would also like to thank Bill Hoffman and Ken Ono for their multiple comments, discussions, and helpful remarks. The authors also thank Wanlin Li who pointed out an error in an earlier draft. Finally, the authors thank the anonymous referee whose comments improved the quality of the exposition. The first author was partially supported by a research assistantship from the Louisiana State University (LSU) Department of Mathematics.
@@ 2. The functions and the arithmetic of elliptic curves
For the hypergeometric datum the function is related to the finite field point counts on the Clausen elliptic curves
Lemma 2.2 (Th. 5 of [Ono98]).
If is prime and then we have
where is the Legendre character on and where
@@
3. Background from Étale Cohomology
In order to prove Theorems 1.2 through 1.9, we describe the traces of an elliptic curve as traces of Frobenius morphisms on an -adic étale sheaf. This perspective allows us to use the resuls and heavy machinery of étale cohomology which we assume the reader is familiar with. In Subsection 3.1, for the reader’s convenience, we recall the necessary background from the theory of étale cohomology. In Subsection 3.2 we recall some background on the sheaves which encode finite-field point counts on families of elliptic curves. Those sheaves “are” essentially the rational -adic Tate modules when the family is viewed as an elliptic curve defined over the function field (see Section 6).
@@ 3. Background from Étale Cohomology
In order to prove Theorems 1.2 and 1.5, we will describe the moments of the values as sums of the form where is an appropriately chosen étale sheaf. To determine the asymptotics of these moments, we express the sums over as sums of traces of acting on cohomology groups associated to the sheaf
@@
[FKMS19, Section 3.4 and Example 4.2.2] Let be a prime a character on and Then there exists a rank- -adic sheaf , pure of weight , lisse on and tamely ramified at and such that
@@
We first recall some elementary properties of
Lemma 3.8 (Section 3 of [Mic95]).
@@
Parts (1) and (2) are Lemma 3.1 of [Mic95]. To see part (3), note that the smooth base change theorem (see Chapter VI, Corollary 4.2 of [Mil80]) implies that is isomorphic as a Galois representation to the where is viewed as an elliptic curve over The self-duality then follows by Poincaré duality (see Chapter VI, Corollary 11.2 of [Mil80]) which in this case is nothing but the Weil pairing. ∎
@@ 4. Background from Representation Theory
Here we recall facts from representation theory which allow for computation of moments using the representations described in Section 3. Aside from the technical details, this section is essentially a generalization of some of the tools used in [OSS23, OPSS24, Gro24]. More precisely, the results in this section combined with Lemma 3.9 elucidate why the Chebyshev polynomial sums are precisely the sums with small asymptotic magnitude. Furthermore, this section elucidates why the Catalan numbers are the moments and transforms the induction arguments on Chebyshev polynomials into direct explicit formulas.
@@
In order to compute the moments in Theorems 1.2 and 1.8, we decompose the tensor product representations into a direct sum of symmetric power representations and a trivial representation. The aim of this is to isolate the irreducible nontrivial components, which by Lemma 3.4, Theorem 3.2, and Theorem 3.3 have small contributions to the moments. The following well-known lemma (for example, see 2.2. of [LZ10]) provides the required direct sum decomposition.
@@
If then this is trivial. Now suppose Using Lemma 4.3, we have that
@@
Putting all this together, we have our claim. ∎
@@
In order to use the geometric monodromy computation of Deligne in Lemma 3.9, we connect the irreducibility of with Recall that the geometric mondromy group is the Zariski closure of in The following well-known lemma (for example, see Exercise 7.1.6 of [Kow14]) relates the irreducibility of a representation
@@
Here we write the proof for the reader’s convenience. If is irreducible, it follows immediately that acts irreducibly. Now, suppose the standard representation of is irreducible, and let be a subspace of such that
@@
It is clear that is a Zariski closed algebraic subgroup of which contains This implies that Since acts irreducibly, or proving our claim. ∎
@@
Let be the basis in the statement of the lemma, and let be an element of If the subspace generated by is invariant under the action of a direct computation shows that ∎
@@ 5. Moments of Frobenius Traces
Here we assemble the results from Sections 3 and 4 to compute the moments and mixed-moments that are required for Theorems 1.2 through 1.9.
@@
where the representations are irreducible pairwise nonisomorphic nontrivial representations of and where is the trivial representation. Part (1) of Lemma 3.9 states that the Zariski closure of Therefore, by Lemma 4.7, we have that the are precisely the symmetric power representations and are the multiplicities in Lemma 4.3. Now, note that by Theorem 3.6, we have that the Zariski closure of is given by Therefore, Lemma 4.8 implies that for we have that is an irreducible representation of It is also clear that and are irreducible. Therefore, the multiplicity of the trivial representation in the decomposition of is equal to and then Lemma 4.3 shows our claim. ∎
@@
Remark 5.4.
@@
by identifying the sum with the trace of Frobenius on the cohomology group where Now, for let and let be the family defined in Section 2. In this case, we have an explicit description of the Frobenius traces on this “mysterious” cohomology group in terms of the traces of Hecke operators on spaces of cusp frorms. For example, if and is even, then (see Section 4.2 of [Gro24]) the trace of the Frobenius map is given by
@@
Now we prove Proposition 5.1.
Proof of Proposition 5.1.
@@
Now, note that these irreducible representations are symmetric powers of By Lemma 4.6, we have that is irreducible. Moreover, it is clear that this representation is never trivial. Therefore, the same argument as in the proof of Lemma 5.3 shows our claim. ∎
@@
which proves our claim. ∎
@@ 6. Tate Modules and Isogenies
First, by Lemma 3.10 and our assumption of good or multiplicative reduction, it is clear that and must have the same points of ramification. By Lemma 4.9, we have that must be lisse on Furthermore, by Lemma 3.9, there exists such that if , then and are tamely ramified at and therefore, is tamely ramified at Therefore, is geometrically constant (see Example 5.2 (f) of Chapter I of [Mil80]). ∎
@@
Now, take to be a nonsquare in and consider By comparing the traces of Frobenius of and the quasi-orthogonality relations of irreducible sheaves (see Theorem 5.2 of [FKMS19]) imply that, up to enlarging we have for all primes ∎
@@ 7. Some Distributions
To obtain Corollaries 1.4 and 1.6, we combine Theorems 1.2 and 1.5 with the method of moments. In this section, we compute the required probability density functions associated to a given sequence of moments or mixed moments.
@@
and our claim follows immediately. ∎
@@
Applying (1) of Lemma 7.3 with and we obtain
The residue theorem then implies that
By Lemma 7.4, our claim follows. ∎
8. Proofs of the Main Results
Proof of Theorem 1.2.
Proof of Theorem 1.5.
Proof of Theorem 1.8.